# Q2E: Query-to-Event Decomposition for Zero-Shot Multilingual Text-to-Video Retrieval

### v1

@@ A.4.5 Full Video Captioner
[…]
[ Video Caption using Frame Captions ] You will receive individual descriptions of consecutive frames from a single video, along with the original transcript of the audio (ASR) from the entire video. Your task is to provide a detailed description of the whole video, summarizing the key actions, participants, and context of the event depicted in the video.
# Instructions:
- Use the original ASR to enhance your understanding of the video context and inform your description of the whole video.
- Focus only on relevant aspects, such as what is happening, who is involved, and the overall context or setting.
- Ignore unrelated visual details like colors, shapes, or other non-essential information.
- Your summary should concisely and accurately capture the essence of the event, integrating insights from the frame descriptions, the Original ASR and the ASR summary.
Only output the video summary, without any additional commentary or content.
# Frame Descriptions:
## Frame [i] Description
{{frame_i_description}}
# Original ASR Transcript (Entire Video):
{{ASR}}
# Summary:
A.4.6 Refining Audio Transcripts
As illustrated in Fig. 2 (left orange box), we utilized an LLM to refine the translations generated by the multilingual-ASR and a translation-specific model. The following prompt was employed to refine the audio transcript using three inputs: (1) the original language transcript from ASR, (2) the translated transcript from the multilingual-ASR, and (3) the translated transcript from the translator.
[…]
[ Refine Audio Transcript ] You are an expert in language translation and text refinement. Your task is to produce a highly accurate and natural-sounding translation by analyzing the original transcript and two provided translations.
# Instructions:
- Combine the strengths of both translations and use the original transcript as a reference to improve accuracy and contextual meaning.
- Prioritize clarity, fluency, and maintaining the original intent and tone of the text.
- Where translations differ significantly, cross-check with the original transcript for fidelity.
- Ensure the final output is free from any awkward phrasing or mistranslation.
- If the transcripts can’t be refined with reasonable accuracy, just output "Not Available".
# Transcripts:
## Original Transcript:
{{original_language_asr}}
## Translation 1:
{{whisper_translated_asr}}
## Translation 2:
{{nllb_translated_asr}}
Output the refined translation only, without additional notes or explanations.
# Refined Translation:
A.5 Use of AI Assistance

### v2

@@ A.6.5 Full Video Captioner
Video Caption using Frame Captions You will receive individual descriptions of consecutive frames from a single video, along with the original transcript of the audio (ASR) from the entire video. Your task is to provide a detailed description of the whole video, summarizing the key actions, participants, and context of the event depicted in the video. # Instructions: - Use the original ASR to enhance your understanding of the video context and inform your description of the whole video. - Focus only on relevant aspects, such as what is happening, who is involved, and the overall context or setting. - Ignore unrelated visual details like colors, shapes, or other non-essential information. - Your summary should concisely and accurately capture the essence of the event, integrating insights from the frame descriptions, the Original ASR and the ASR summary. Only output the video summary, without any additional commentary or content. # Frame Descriptions: ## Frame [i] Description {{frame_i_description}} # Original ASR Transcript (Entire Video): {{ASR}} # Summary:
A.6.6 Refining Audio Transcripts
As illustrated in Fig.˜2 (left orange box), we utilized an LLM to refine the translations generated by the multilingual-ASR and a translation-specific model. The following prompt was employed to refine the audio transcript using three inputs: (1) the original language transcript from ASR, (2) the translated transcript from the multilingual-ASR, and (3) the translated transcript from the translator.
Refine Audio Transcript You are an expert in language translation and text refinement. Your task is to produce a highly accurate and natural-sounding translation by analyzing the original transcript and two provided translations. # Instructions: - Combine the strengths of both translations and use the original transcript as a reference to improve accuracy and contextual meaning. - Prioritize clarity, fluency, and maintaining the original intent and tone of the text. - Where translations differ significantly, cross-check with the original transcript for fidelity. - Ensure the final output is free from any awkward phrasing or mistranslation. - If the transcripts can’t be refined with reasonable accuracy, just output "Not Available". # Transcripts: ## Original Transcript: {{original_language_asr}} ## Translation 1: {{whisper_translated_asr}} ## Translation 2: {{nllb_translated_asr}} Output the refined translation only, without additional notes or explanations. # Refined Translation:
A.7 Use of AI Assistance